A 800N man is playing on a seesaw with son who has a weight of 200N. If the child is sitting 3 meters away, how far will the man have to sit to balance the seesaw?
1
Expert's answer
2020-12-02T07:14:37-0500

The clockwise moment has to be equal to the anticlockwise moment in order for the seesaw to be balanced.

"Moment=Force \times perpendicular \hspace0.05cm 
distance\hspace0.05cm from\hspace0.05cm the\hspace0.05cm centre."

"\therefore" The moment of the man = the moment of the boy

"force(man)= 800N \\
force(boy)= 200N\\
Perpendicular\hspace0.05cm dist.(man)= ?\\
Perpendicular\hspace0.05cmdist.(boy)= 3\hspace0.05cmmetres\\"

"F_{m}\times d_{m}=F_{b}\times d_{b}\\
800\times d_{m}= 200\times 3\\
d_{m}= \frac{800}{(200\times 3)} metres \\
d_{m}= 1.33\hspace0.05cm metres"

The man would have to sit 1.333 metres away